News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Monday 15 May 2000 7.30am EDT

During the last week, the daily load was re-built to continue operations without use of the HETG to allow time for proper testing of B-side grating operations. A real-time procedure was run at 5.30pm on 5/14 to verify the operation of both the HETG and LETG using the B-side equipment (MCE-B). Both gratings were inserted and retracted using both hardware commands and software commands (the latter as used from the daily load). All tests were successful and is was agreed that normal operations using the B-side could begin this week.

The observation schedule from May 6 through the coming week is given below together with annotations describing the HETG anomaly and recovery to B-side operation.

------------------------------------------
XTE J1550-564          ACIS-S/HETG     May  6  HETG retracted: no switch trip
Radiation Belts
SN 1993J               ACIS-S          May  7
M82                    ACIS-I
Proxima Centauri       ACIS-S/HETG             HETG failed insert
Abell 1995             ACIS-S          May  8
XTE J1550-564          ACIS-S/HETG             observation made w/o HETG
Proxima Centauri       ACIS-S/HETG             observation made w/o HETG
Radiation Belts                        May  9
ZW 3146                ACIS-I                  start of rebuilt load
NGC 6543               ACIS-S          May 10
RX J2247+0337          ACIS-I          May 11
Radiation Belts
0509-67.5              ACIS-S          May 12
NGC 4736               ACIS-S
A2199                  ACIS-S          May 13
MS 2053.7-0449         ACIS-I
Radiation Belts                        May 14
1708-408               ACIS-S
NGC 6752               ACIS-S          May 15
Q0059-2735             ACIS-S
1E 0035.4-7230         ACIS-S
Abell 3128             ACIS-I
Rho Oph A              ACIS-I
Radiation Belts                        May 16
RX J0439.8-6809        ACIS-S          May 17
NGC 5128               ACIS-I
EX HYA                 ACIS-S/HETG     May 18  HETG on B-side
BAL 1312+2319          ACIS-S
MS 1455.0+2232         ACIS-I          May 19
NGC 4374 (M84)         ACIS-S
CAS A                  ACIS-I
Radiation Belts
BAL 1556+3517          ACIS-S          May 20
NGC 4244               ACIS-S
Cygnus A               ACIS-S
CAS A                  ACIS-S          May 21
4U 0142+61             ACIS-S
Cygnus Loop SW         ACIS-I

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports